OpenStudy (summersnow8):

Find the mechanical advantage of a pipe vise if it has a handle diameter of 16 in., a screw diameter of 0.75 in., and a screw pitch of 0.125 in. FR * pitch = FE * 2pi * r MA = 2 pi r / pitch = 2 pi (d/2) / pitch = pi d / pitch = pi * 16 in / 0.125 in = 128 pi = 804.248 = 804 how many significant figures should my answer be? 2 or 3?

OpenStudy (michele_laino):

your answer is correct! I think that better is to round off to the nearest hundred, namely: MA= 800

OpenStudy (summersnow8):

actually, the answer should be 400... but I am not sure how to get that answer

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (michele_laino):

maybe we have to consider the efficiency of the screw which can be 0.5

OpenStudy (michele_laino):

if we have to consider the efficiency of the screw, then the requested advantage is: \[\Large MA = \frac{{\pi D}}{{pitch}} \times \eta \] where \eta is the efficiency of the screw Please keep in mind that a screw is a simple machine, so it has an own efficiency

OpenStudy (michele_laino):

of course, we have: \[\Large \eta = 50\% \]

OpenStudy (summersnow8):

@Michele_Laino MA = (π * 16 in / 2) * 50% = 12? how do I get 400?

OpenStudy (michele_laino):

Sorry, I got this: \[MA = \frac{{3.14 \times 16}}{{0.125}} = 401.92\]

OpenStudy (summersnow8):

why do we use 16 in and not .75 in ?

OpenStudy (summersnow8):

@Michele_Laino